We considered the differences between the temperature readings in January 1 of 1968 and 2008 at 51 locations in the continental US in Exercise 5.19. The mean and standard deviation of the reported differences are 1.1 degrees and 4.9 degrees respectively. 

(a) Calculate a 90% confidence interval for the average difference between the temperature measurements between 1968 and 2008. 
Answer #1

Confidence interval formula for paired t test.
✓ Et*

Sample size = 51

Degree of freedom = n - 1 = 51 - 1 = 50

t critical value is = 1.68.......................by using t table or by using Excel command =TINV(0.1,50)

Substitute here known values ,

* 89 TETI<=

=>(-0.05, 2.25)

Lower bound = -0.05

Upper bound = 2.25